Find the using variation general solution of parameters y^y-2y=2et Y
Answer #1

We can solve the given initial value problem by using method of variation of parameters method.

So the most general solution is   Y = YC + YP

YC = Homogeneous solutin

YP = Particular solution.

Now the given initil value problem is y’’ - y’ - 2y = 2e-t

So let us take the homogeneous equation for solving

So y’’ - y’ - 2y = 0    -------------------(4)

Now let us find the characteristic equation for homogeneous equation by assuming the solution y = ert not equla to Zero.

So y’ = r ert

And y’’ = r2ert

so substitute in (1)

r2ert - rert - 2ert = 0

ert (r2 - r - 2) = 0

ert is not equal to zero so r2 - r - 2 = 0

so r2 - r - 2 = 0

r2 + r - 2r - 2 = 0

r (r+ 1) - 2(r + 1) = 0

(r - 2) (r + 1) = 0

r1 = 2 and r2 = -1

so when the characteristic equation roots are real and differrent then homogeneous solution is in the form as given below.

Yc = C1er1t + C2er2t

Yc = C1e2t + C2e-t

So now we need to find the wronskian for particular solution

Yc = C1y1(t) + C2y2(t)

So y1(t) = e2t and y2(t) = e-t

y1’(t) = 2e2t and y2’(t) = -e-t

And  W(y1 , y2) = y1(t) y2’(t) - y1’(t) y2(t)

W(y1 , y2) = (e2t) (-e-t) – (2e2t) (e-t)

W(y1 , y2) = -(e2t-t) – (2e2t-t)

W(y1 , y2) = -(et) – (2et)

W(y1 , y2) =  – (3et)

And non homogeneous term g(t) = 2e-t


How the particular solution Up = my, (4) f y lt.gif wly, Y₂) .dt Yalt). g(t) -dt & Yalt) wly, Y2) et rzet) .de -Bet Ypa -e26